OpenStudy (anonymous):

All together, Dion and Trevor worked a total of 38 hours last week. If Dion worked d hours, which expression represents the number of hours Trevor worked?

OpenStudy (godlovesme):

options?

OpenStudy (anonymous):

mc010-1.jpg mc010-2.jpg mc010-3.jpg mc010-4.jpg

OpenStudy (anonymous):

38 + D 38 - d 38/d 38+d ------- 2

OpenStudy (mathstudent55):

Let's say Dion worked d hours, and Trevor worked t hours. Then the total both worked is d + t We know the total they worked is 38, so we know d + t = 38 Now we want to know how many hours Trevor worked, so we need to find what t is equal to. We solve d + t = 38 for t. We subtract d from both sides to isolate t, and we get: d + t = 38 -d - d ---------------- t = 38 - d